Find the radius of a spherical conductor so has to have a capacitance of 10 micro farad ?
 

Grade:12

1 Answers

Arun
5 years ago

Capacitance of a spherical conductor = 4πϵ0a F, where

ϵ0 = permittivity of free space = 8.85418782 × 10^(-12)

a = radius of the sphere in m

If C is 10μF = 10^(-5) F, then,

a = 10^(-5)/(4πx8.85418782 × 10^(-12) = 89875.52 m